# Exploring Young Stellar Variability with Gaia DR3 light curves

### v1

@@ 6.2.5 Bursters
B-variables were previously associated with high levels of accretion based on correlations with various accretion proxies, from UV excess to variability in colour behaviour and H (Stauffer et al., 2014; Cody et al., 2017; Venuti et al., 2021; Hillenbrand et al., 2022). In our sample, this association with star-disk-interaction phenomenology is only supported for sources in our primary B sample, for which the inner disk and CTTS frequency were 45.18% and 45.49%, while for the secondary sample these are 12.22 and 7.04% respectively.
As defined by Cody et al. (2014), the B class might constitute a family of bursting behaviours, rather than relating to a single physical mechanism. In a study focused solely on B-variability, Cody et al. (2017) finds varied B behaviour, with some sources showing a continuous series of bursts, while others presenting short-lived burst episodes on top of underlying irregular variability (see also Venuti et al., 2021). This family of prototypical B behaviour is verified among our primary B sample, and illustrated in Fig. 19, where we show examples of isolated bursts, including both short-lived (GDR3 510735868513196160) and longer-lived bursts (GDR3 6236134222781410944). Fig. 19 also includes an example of a quasi-periodic B variable with intermediate timescale, GDR3 2069894002944444928, which may be related to the phenomenology behind the DQ Tau system (Tofflemire et al., 2017), where two stars orbit each other in a highly eccentric orbit and their magnetic interaction with circumbinary and circumstellar material yields enhanced accretion near periastron, resulting in the observation of regular bursts.
Following the discussion on outlier removal in Sect. 3.4 and App. A.3, we note that Gaia’s sparse cadence is expected to yield an under-detection of variables with short-lived bursts. As reported by Cody et al. (2017), only 1 burst out of 29 monitored with high-cadence K2 light curves showed burst durations longer than 10 days. If observed by Gaia, such a sample would likely have burst events ruled out as outliers, unless the bursting feature coincides with a cusp, as in the case of a short-lived burst variable GDR3 6236134222781410944 shown in Fig. 19.
Finally, we did not find examples of extreme burst events, such as a FUor or EXors, among our sample of B sources. However, we verified the existence of some sharp long-term bursts among our long-trend variables (see the top panel of Fig. 22).
@@ 6.2.6 Stochastics
S variables (Fig. 20), are dominated by irregular variability without detectable periodicity or a clear preference for bursting or dimming behaviour. Fig. 14 shows that even our primary sample shows systematically lower amplitudes compared to previous studies. Previous high-cadence surveys associated S-variables with a combination of the erratic routine variability introduced by accretion combined with viewing angles close to face-on, which hindered the rotational modulation of accretion-induced variability (Stauffer et al., 2016). This scenario is supported by our main S-sample, which has inner disk and CTTS frequencies of 43.26 and 49.04%, respectively, along 2.26 times as many intermediate timescales as rotational ones. However, visual inspection of S light curves shows that a fraction of those are multi-mode variables with alternating bursting and dipping behaviour across the observed time (see Fig. 20 bottom panel). In contrast, our secondary sample has very low inner disk and CTTS frequency (18, and 7.36%), with a wide spread of timescales and no clear preference for rotational or intermediate timescales. However, when disaggregated by amplitude, this sample reveals an excess of sources with low-amplitude (below 0.1 mag) fast rotation timescale (below 5 days), which may suggest a link between this sample and cool spot dominated light curves close to Gaia’s detection limit.
@@ 6.2.7 Eclipsing binaries
The EB class in Cody et al. (2014) targeted prototypical eclipsing binary light curves with strictly periodic asymmetric light curves. We found 129 EB sources. A search for counterparts within the Simbad Database (Wenger et al., 2000) reveals that only 4 of our EB sources were previously identified as such in the literature. We note that although sources classified as EB generally followed the prototype of strictly periodic eclipses, our automated method for timescale derivation with the GLS (Sect. 4.2.1) tends to detect harmonics of the true period as the best timescale. We show two examples of EB light curves in Fig. 21, where in both cases the GLS returned half of the true period as the best timescale. In the case of the source GDR3 5414630399235694976 (top plots), secondary and primary eclipses are relatively similar in depth with the waveform estimated based on the biased timescale representing the true waveform relatively well and yielding a very low Q-value. However, in cases like the source GDR3 2019728853698920704 (bottom plots), as the secondary eclipse is much shallower, waveform estimation based on the harmonic of the true timescale inflates Q values, resulting in the miss-classification of the source as a QPD.
6.2.8 Long-trends
Although the L-variables do not have a Q-index (because the timescale was too large to allow a relevant waveform to be derived from the phase-folded curve), we still computed their M-index. In the primary sample, 21% Ls have bursting behaviour and 30% have dipping behaviour. For the secondary sample, 21% Ls have bursting behaviour, and 26% have dipping behaviour. Ls tend to have larger amplitudes than the other classes, which is expected since there is an empirical correlation between the timescale of an event and its amplitude (See Fig. 3 of Fischer et al., 2023).
@@ 6.3 Comparison across surveys
To further investigate the consistency of Q and M indices across different surveys, we cross-matched our variable sample with previous Q&M studies using K2 (Cody et al., 2022) and ZTF (Hillenbrand et al., 2022; Wang et al., 2023; Jiang & Hillenbrand, 2024). Figs. 23 and 24 show examples of light curves of sources with observations in multiple surveys. We emphasise that none of the light curves in question included common seasons of observations.
@@ 6.3.1 High-cadence space-based surveys
We found only 13 sources in common between our variable sample and the sources observed by Cody et al. (2022) with K2 in Taurus (see Fig. 23). Hence, we limit our comparison to a qualitative discussion. Since K2 and GDR3 have very different cadences and timespans, our comparison focuses on their M-index behaviour, where classification was considered equivalent if the source was detected as dipping, symmetric or bursting in both surveys. Although variability amplitudes were generally consistent across surveys, Q&M-indices often had large enough differences to produce distinct morphological classes.
Of the three symmetric sources in K2, only one was also symmetric in GDR3, while the others were dipping. However, since these surveys cover different seasons of observation, it is impossible to infer whether discrepant classifications result from a change in dominant variability mode or the result of different cadences of observations. Three B-variables in K2 were also B in GDR3. One of these sources is GDR3 3409319239697943296 (Fig. 23), which, although bursting in both surveys, was quasiperiodic in our study and aperiodic when observed by K2. Of the four dipping variables in K2, three were also dipping in GDR3. Cody et al. (2022) searched for multiple variability modes by computing Q&M twice: first on the original light curve and then on the residual data. Two out of three multimode sources in common with GDR3 have one of their morphological classes matched to GDR3. An example is shown in Fig. 23, where the QPD source GDR3 3406128761895775872 was attributed to the QPS and APD classes in the K2 analysis.
@@ 6.3.2 Comparison with ZTF
ZTF light curves have a similar duration to GDR3, but are noisier due to ground-based observations. We found 165 sources in common with ZTF studies. Of these, 25 sources were sorted as L by our method, thus having no Q-index. Two example light curves are displayed in Fig. 24. GDR3 2067026678480142464, on the top panel, shows a similar behaviour in both surveys. Although it was classified as a long-trend by us, its M-index implies bursting behaviour. GDR3 M-index is slightly shifted towards more symmetric compared to ZTF. GDR3 216701033028221312 (bottom panel) has a similar behaviour in both surveys, with matching timescales. However, the dispersion is considerably more significant in the ZTF data than in GDR3, resulting in a higher Q-index value for ZTF. Additionally, the ZTF light curve contains significantly more data points, contributing to its more symmetric behaviour compared to GDR3.
Fig. 25 shows the Q-M diagrams for each of the GDR3 classes (except L), with both the GDR3 and ZTF results scattered. At a population level, results between the two surveys are relatively consistent, especially in terms of light-curve symmetry. There were very few cases where variables were detected in radically different behaviours, for example, switching from bursts to dips. For periodic and quasi-periodic sources, we find that differences in derived timescales were either due to harmonics, or could be explained by the 1-day aliasing caused by ZTF’s ground-based observations. For aperiodic sources, the timescales were significantly different. However, these can be traced back to different methodologies used for the aperiodic timescale derivation. Given the different Q-index thresholds adopted in ZTF surveys, most QPS variables in Gaia were also QPS in ZTF. A few dippers in ZTF were detected as QPS or S-variables in Gaia, although we note that these are generally close to the M-threshold in our study. Most dippers (QPDs or APDs) in Gaia were also dippers in ZTF. Although a fraction of Gaia’s B-variables are recovered by ZTF, most bursts in Gaia are detected as symmetric in ZTF. Nevertheless, we note that because of their lower cadence, both Gaia and ZTF are expected to be inefficient in detecting bursts.
@@ 7 Summary and conclusions
In this paper, we presented the results of a pilot study employing the asymmetry and periodicity variability metrics to characterise an all-sky sample of YSO candidates based on their G-band photometric variability observed with Gaia. We have adapted the methodology behind the Q&M-indices for application to Gaia’s long-term sparse light curves. This methodology was further developed into an automated workflow for managing large datasets. Starting from the sample of 79 375 variables identified in GDR3 as YSO candidates based on their variability, we pre-processed sources to filter for a low number of epochs, remove contaminants and identify sources likely affected by spurious instrumental variability. Based on a sample of 28 179 sources, we inferred characteristic timescales of variability considering both the possibility of periodic and aperiodic variability. The timescale distribution revealed a bimodality of rotational and intermediate timescales, with the former showing slightly lower variability amplitudes on average than the latter. Rotational timescales were more frequent for sources with enhanced periodicity and intermediate timescales for aperiodic sources. We could successfully derive Q&M indices for 23 213 sources. We complemented our analysis with observations of H as a proxy of ongoing accretion and as a proxy of the presence of circumstellar material and the evolution stage of YSOs. We found that WTTS diskless sources tend to show more periodic and symmetric variability, while CTTS and WTTS disk-bearing sources tend to be more asymmetric and less periodic.
@@
We used the derived Q&M indices to sort sources into one of six variability morphology classes: Strictly Periodic, Quasiperiodic Symmetric, Stochastic, Quasiperiodic Dipper, Aperiodic Dipper, Burst, or Eclipsing binary. 4 966 additional sources had detected timescales longer than half of the total observed time and were classified as Long-trend, with only M-index derived. The subsamples in each of these classes were further examined in terms of Q&M-indices, variability amplitudes, timescales, and accretion and disk proxies with possible links to well-known variability mechanisms discussed. This examination revealed a wider variety of variability behaviour than previous studies, which we attribute to Gaia’s longer timespan. We found sources displaying the usual variability caused by rotational modulation by magnetically induced spots, or accretion induced hot spots, and both bursting and dipping behaviour at rotational timescales, likely linked to star-disk-interaction. Moreover, we found a relevant fraction of sources showing dipping and bursting behaviour at intermediate timescale, which we tentatively linked to the phenomenology behind RW Aur fades, and DQ Tau-type bursting variability.
@@
We qualitatively compared our Q&M-based results to previous studies, considering both the results of high-cadence surveys (CoRoT and K2) and ground-based surveys (ZTF) and including discussion at the individual and population levels. Although large survey-to-survey differences are found, M-index trends seem generally consistent. When considering survey-specific biases (caused by cadence, timespan, precision and age), we found that the differences in the incidence of morphological classes across surveys can be explained. We note a lack of available observed datasets with common sources observed by Gaia and high-cadence surveys, which would allow us to probe the appropriateness of Q&M thresholds for classification across surveys. Future studies should consider assessing the influence of survey-to-survey variations on these thresholds based on simulated data. Finally, Gaia DR4 will allow us to expand these results to an even broader all-sky coverage, thanks to the longer timespan of observations and improved calibration.

### v2

@@ 6.2.5 Bursters
Based on correlations with various accretion proxies, from UV excess to variability in colour behaviour and H, B variables were previously associated with high levels of accretion (Stauffer2014AJ....147...83S; Cody2017; Venuti_2021; Hillenbrand2022). In our sample, this association with star-disc-interaction phenomenology is only supported for sources in our primary sample, for which the inner disc and CTTS frequency were 45.18% and 45.49%, while for the secondary sample these are 12.22 and 7.04% respectively. Among these with evidence of a disc or ongoing accretion, we note that the earlier and later type sources share statistically different distributions of amplitude, with the distribution of earlier types (mean: mag) shifted towards higher amplitudes in comparison with later type sources ( mag).
As defined by Cody_14, the B class might constitute a family of bursting behaviours, rather than relating to a single physical mechanism. In a study focused solely on B variables, Cody2017 finds varied bursting behaviour, with some sources showing a continuous series of bursts, while others presenting short-lived burst episodes on top of underlying irregular variability (see also Venuti_2021). This family of bursting behaviours is verified among our primary sample, and illustrated in Fig. 21, where we show examples of isolated bursts, including both short-lived (GDR3 510735868513196160) and longer-lived bursts (GDR3 6236134222781410944). Fig. 21 also includes an example of a quasi-periodic B variable with intermediate timescale, GDR3 2069894002944444928, which may be related to the phenomenology behind the DQ Tau system (Tofflemire2017ApJ...835....8T), where two stars orbit each other in a highly eccentric orbit and their magnetic interaction with circumbinary and circumstellar material yields enhanced accretion near periastron, resulting in the observation of regular bursts. Following the discussion on outlier removal in Sect. 3.4 and App. A.4, we note that the sparse cadence of Gaia is expected to yield an under-detection of variables with short-lived bursts. As reported by Cody2017, only 1 burst out of 29 monitored with high-cadence K2 light curves showed burst durations longer than 10 days. If observed by Gaia, such a sample would likely have burst events ruled out as outliers, unless the bursting feature coincides with a cusp, as in the case of a short-lived burst variable GDR3 6236134222781410944 shown in Fig. 21. Finally, we did not find examples of extreme burst events, such as a FUor or EXors. However, we verified the existence of some sharp long-term bursts among our long-trend variables (e. g., Fig. 24).
@@ 6.2.6 Stochastics
The light curves of S variables (Fig. 22) are dominated by irregular variability without detectable periodicity or a clear preference for bursting or dimming behaviour. Fig. 16 shows that even our primary sample shows systematically lower amplitudes compared to previous studies. Previous high-cadence surveys associated S-variables with a combination of the erratic routine variability introduced by accretion, combined with viewing angles close to face-on, which hindered the rotational modulation of accretion-induced variability (Stauffer2016AJ....151...60S). This scenario is supported by our primary S sample, which has inner disc and CTTS frequencies of 43.26 and 49.04%, respectively, along with 2.3 times as many intermediate timescales as rotational ones. However, visual inspection of S light curves shows that a fraction of those are multi-mode variables with alternating bursting and dipping behaviour across the observed time (see Fig. 22 bottom panel). In contrast, our secondary sample has very low inner disc and CTTS frequency (18 and 7.36%), with a wide spread of timescales and no clear preference for rotational or intermediate timescales. However, when disaggregated by amplitude, this sample reveals an excess of sources with low-amplitude (below 0.1 mag) fast rotation timescales (below 5 days), which may suggest a link between this sample and cool spot-dominated light curves close to the Gaia detection limit.
@@ 6.2.7 Eclipsing binaries
The EB class in Cody_14 targeted prototypical eclipsing binary light curves with strictly periodic asymmetric light curves. We found 129 EB sources. A search for counterparts in the Simbad Database (2000A&AS..143....9W) reveals that only 4 of our EB sources were previously identified as such in the literature. Although sources classified as EB generally followed the prototype of strictly periodic eclipses, our automated method for timescale derivation using the GLS (Sect. 4.2.1) often detects harmonics of the true period as the best timescale. We show two examples of EB light curves in Fig. 23, where in both cases the GLS returned half of the true period as the best timescale. In the case of the source GDR3 5414630399235694976 (top plots), secondary and primary eclipses are relatively similar in depth, with the waveform estimated based on the biased timescale representing the true waveform relatively well and yielding a very low Q value. However, in cases like the source GDR3 2019728853698920704 (bottom plots), as the secondary eclipse is much shallower, waveform estimation based on the harmonic of the true timescale inflates Q values, resulting in the miss-classification of the source as a QPD.
6.2.8 Long trends
With timescales long for a relevant waveform to be derived from the phase-folded curve, L variables lack a Q index, but we still computed their M index. In the primary sample, 21% Ls have bursting behaviour and 30% have dipping behaviour. For the secondary sample, 21% Ls have bursting behaviour, and 26% have dipping behaviour. L variables tend to have larger amplitudes than the other classes, which is expected since there is an empirical correlation between the timescale of an event and its amplitude (See Fig. 3 of Fischer23). Although we do not verify any preference for bursting or dipping as a function of their spectral types, we find that earlier type sources have typically much lower amplitudes (mean: mag) compared with later type sources ( mag).
@@ 6.3 Comparison across surveys
To further investigate the consistency of Q and M indices across different surveys, we cross-matched our variable sample with previous Q-M studies using K2 (Cody_2022) and ZTF (Hillenbrand2022; Wang2023; JiangHillenbrand2024). Figs. 25 and 26 show examples of light curves of sources with observations in multiple surveys. Q-M indices from the literature are adopted as reported in the source papers. None of these surveys included seasons of observations in common.
@@ 6.3.1 High-cadence space-based surveys
We found only 13 sources in common between our variable sample and the sources observed by Cody_2022 with K2 in Taurus (see Fig. 25). Hence, we limit our comparison to a qualitative discussion. Since K2 and GDR3 have very different cadences and time spans, our comparison focuses on their M index behaviour, where classification was considered equivalent if the source was detected as dipping, symmetric or bursting in both surveys. Although variability amplitudes were generally consistent across surveys, Q-M indices often had large enough differences to produce distinct morphological classes.
Of the three symmetric sources in K2, only one was also symmetric in GDR3, while the others were dipping. However, since these surveys cover different seasons of observation, it is impossible to infer whether discrepant classifications result from a change in dominant variability mode or the result of different cadences of observations. Three B-variables in K2 were also B in GDR3. One of these sources is GDR3 3409319239697943296 (Fig. 25), which, although bursting in both surveys, was quasi-periodic in our study and aperiodic when observed by K2. Of the four dipping variables in K2, three were also dipping in GDR3. Cody_2022 searched for multiple variability modes by computing Q-M twice: first on the original light curve and then on the residual data. Two out of three multi-mode sources in common with GDR3 have one of their morphological classes matched to GDR3. An example is shown in Fig. 25, where the QPD source GDR3 3406128761895775872 was attributed to the QPS and APD classes in the K2 analysis.
@@ 6.3.2 Comparison with ZTF
Light curves observed by ZTF have roughly the same duration as GDR3 but are noisier due to ground-based observations. With 165 sources in common, Fig. 27 compares the Q-M distribution between surveys, with 25 sources omitted as they were classified as L variables (no Q index) in the present survey. At the population level, the two surveys yield broadly consistent results, particularly for light curve symmetry with very few cases where variables were detected in radically different behaviours, for example, switching from bursts to dips. For periodic and quasi-periodic sources, we find that differences in derived timescales were either due to harmonics or to one-day aliasing caused by ZTF ground-based observations. For aperiodic sources, timescales differed substantially, but this was explained by the distinct methods used to derive them. Given the different Q index thresholds adopted in ZTF surveys, most QPS variables in Gaia were also QPS in ZTF. Some dippers in ZTF were identified as QPS or S variables in GDR3, although they are generally close to the M threshold in our study. Most dippers (QPDs or APDs) in Gaia were also dippers in ZTF. Most bursts in Gaia are detected as symmetric in ZTF, although we note that both surveys’ low cadence makes them inefficient in detecting bursts. Fig. 26 shows two examples of sources in common in both surveys. GDR3 2067026678480142464 (top) shows similar behaviour in both surveys, and although we classified it as an L variable, its M index suggests the same bursting behaviour, with a slight shift towards higher symmetry in GDR3 than in ZTF. GDR3 216701033028221312 (bottom) has a similar behaviour in both surveys, with matching timescales but higher phase dispersion in ZTF than in GDR3, resulting in a higher Q index value for ZTF. Additionally, the ZTF light curve contains significantly more data points, contributing to its more symmetric behaviour compared to GDR3.
@@ 7 Summary and conclusions
We presented the results of a pilot study employing the asymmetry and periodicity variability metrics to characterise an all-sky sample of YSO candidates based on their G-band photometric variability observed with Gaia. We adapted the method behind the Q-M indices for application to Gaia’s long-term but sparse light curves. Starting from the sample of 79 375 variables identified in GDR3 as YSO candidates based on their variability, we pre-processed sources to filter for a low number of epochs, remove contaminants, and identify sources that were likely affected by spurious instrumental variability. Based on a sample of 28 179 sources, we inferred characteristic variability timescales by considering the possibility of periodic and aperiodic variability. The timescale distribution revealed a bimodality with rotational and intermediate timescales, with the former showing, on average, lower variability amplitudes than the latter. Rotational timescales were more frequent for sources with a higher periodicity and intermediate timescales for aperiodic sources. We successfully derived Q-M indices for 23 213 sources. We complemented our analysis with observations of H as a proxy of ongoing accretion and as a proxy of the presence of circumstellar material and the evolution stage of YSOs. We found that WTTS discless sources tend to show more periodic and symmetric variability, while CTTS and WTTS disc-bearing sources tend to be more asymmetric and less periodic. We also examined the results as a function of the estimated spectral types for the sample and found that QPS variability is much more common among M type stars, while the incidence of L variability increases towards earlier types.
@@
We used the derived Q-M indices to sort sources into one of six variability morphology classes: Strictly periodic, quasi-periodic symmetric, stochastic, quasi-periodic dipper, aperiodic dipper, burst, or eclipsing binary. The timescales of 4 966 additional sources were longer than half of the total observed time, and these were classified as long trend variables with only their M index derived. The subsamples in each of these classes were further examined in terms of Q-M indices, variability amplitudes, timescales, and accretion and disc proxies, and we discussed possible links to well-known variability mechanisms. This examination revealed a wider variety of variability behaviour than previous studies, which we attribute to the longer time span of Gaia. We found sources displaying the usual variability caused by rotational modulation by magnetically induced spots or accretion-induced hot spots, and both bursting and dipping behaviour at rotational timescales, likely linked to star-disc interaction. Moreover, we found a relevant fraction of sources showing dipping and bursting behaviour at an intermediate timescale, which we tentatively linked to the phenomenology behind RW Aur fades and DQ Tau-type bursting variability.
@@
We qualitatively compared our Q-M-based results with previous studies, including both high-cadence (CoRoT and K2) and ground-based surveys (ZTF), and discussed them at the individual and population levels. Although large survey-to-survey differences were found, the M index trends seem generally consistent. When considering survey-specific biases (caused by cadence, time span, precision, and age), we found that differences in the incidence of morphological classes across surveys can be explained. We note a lack of available observed datasets with common sources observed by Gaia and high-cadence surveys, which would allow us to probe the appropriateness of Q-M thresholds for a classification across surveys. Future studies should consider assessing the effect of survey-to-survey variations on these thresholds based on simulated data. Finally, Gaia DR4 will allow us to expand these results to an even broader all-sky coverage because the time span of the observations is longer and the calibration is improved.